OpenStudy (anonymous):

My answer is C What is the sum of a 30-term arithmetic sequence where the first term is 74 and the last term is -100? (3 points) Answer -468 -442 -416 -390

hartnn (hartnn):

no, its not C, which formula did you use ?

hartnn (hartnn):

you should have used, \(\huge S_n = (n/2)(a_1+a_n)\)

OpenStudy (anonymous):

arithmetic series formula: S(n) = n[a(1) + a(n)]/2

hartnn (hartnn):

correct, so n= 30, a1 = 74, an=-100 what you get as Sn ?

OpenStudy (anonymous):

30

hartnn (hartnn):

:O

hartnn (hartnn):

which calculator are you using ? make sure you are calculating (30/2)(74-100)=...?

OpenStudy (anonymous):

30 (74 - 100)/2

hartnn (hartnn):

which equals ?

OpenStudy (anonymous):

wait why does ours look different?

OpenStudy (anonymous):

1st term minus last term i thought

hartnn (hartnn):

its same, it just 'looks' different,30 (74 - 100)/2 = (30/2)(74-100)

OpenStudy (anonymous):

thank god

hartnn (hartnn):

its 1st term + last term. a1+an

OpenStudy (anonymous):

-390

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

wow thank you:)

hartnn (hartnn):

\[ \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{^}}\color{#e6e600}{\text{_}}\color{green}{\text{^}}\color{blue}{\text{}}\end{array} \]